OpenStudy (anonymous):

Angle θ lies in the fourth quadrant, and cot θ =-8/15 .

OpenStudy (anonymous):

sin(theta)=? cos(theta)=?

OpenStudy (anonymous):

\[\cot \theta=\frac{ \cos \theta }{ \sin \theta}\]